9.As a quality test of ball bearings, you drop bearings (small metal balls), with zero initial velocity, from a height Hi = 12.18 m onto a thick steel plate. The bearings bounce back up to a final height Hf = 10.17 m.

If the bearings are in contact with the steel plate for a time interval of 0.010 s, what is the magnitude of the average acceleration of the bearings while they are in contact with the plate?

Give your answer in m/s2 to at least 2 decimal places.
Answer #1

Given that

initial height hi=12.18 m

finial height hf=10.17 m

time interval t=0.01 sec

now we find the average acceleration

avergae acceleration a=2(hf-hi)/t^2=2(10.17-12.18)/0.01^2

=-40.2*10^3 m/s^2